OpenStudy (priyar):

Let f:(-1,1) -> B, be a function defined by f(x)=tan inverse(2x/1-x^2), then f is both one-one and onto when B is the interval:

OpenStudy (priyar):

@pooja195 @ParthKohli

OpenStudy (priyar):

@roadjester

OpenStudy (anonymous):

is your B: \(\left(-\frac{\pi}{2}, \frac{\pi}{2} \right)\)

OpenStudy (priyar):

yes...but can u tell me how u arrived at the answer?

OpenStudy (mathmale):

One has to read your post several times in an effort to determine what you're asking. Could you possibly be more specific in stating the goal of this problem? On the side: Why not graph the tangent function and the inverse tangent function, as a review? Specify the domain and range of each.

OpenStudy (priyar):

r u asking the goal of the q? its simple its to find the range of the given function

OpenStudy (priyar):

ehich part u didn't understand? @mathmale ?

OpenStudy (priyar):

*which

OpenStudy (priyar):

and i can't plot a graph coz its an objective q..

OpenStudy (priyar):

i got the answer!